OpenStudy (anonymous):

The Orlando Science Center wants to frame a rectangular display that is 72 inches long and 36 inches wide. How many inches of framing material is needed to surround the display?

OpenStudy (jwheele1):

it is asking for perimeter

OpenStudy (jwheele1):

of a rectangle. do you know the formula for that?

OpenStudy (anonymous):

2l X 2W

OpenStudy (jwheele1):

2L PLUS 2W

OpenStudy (jwheele1):

or side +side +side +side

OpenStudy (anonymous):

ohhh kk

OpenStudy (jwheele1):

so, 2L + 2W 2(72) + 2 (36) = ?

OpenStudy (anonymous):

5256

OpenStudy (jwheele1):

lets check that. 2 * 72 = 144 2*36 = 72 --------- 144 + 72= 216

OpenStudy (jwheele1):

216 inches

OpenStudy (jwheele1):

BUT....how did you get your answer

OpenStudy (jwheele1):

order of operations states that you must do all multiplication before addition

OpenStudy (anonymous):

ohhhhhhhhhhhhhh

OpenStudy (jwheele1):

PEMDAS Please Excuse My Dear Aunt Sally P is for parenthesis first...do whats inside them first E is for exponents next. M and D are together and they stand for multiply and divide...you do these before addition and subtraction and you do these from left to right. A is for addition and S is for subtraction. You should do these from left to right. for good habit.

OpenStudy (anonymous):

ty

OpenStudy (jwheele1):

yw